A heater element has a voltage rating of 230 V. If this heater element converts 2.5kWh in a span of 1.5hrs, determine the current flowing in the heater element during this period
1
Expert's answer
2021-09-13T16:52:36-0400

Power dissipated by the heater is

"\dfrac{2.5kWh}{1.5h}=1.667kW"

"Power=Voltage \times{Current}\\
Current=\dfrac{Power}{Voltage}=\dfrac{1.667\times1000}{230}=7.2A"

Current is therefore 7.2A